OpenStudy (dcolley):

Find the quadratic function y=ax^2+bx+c whose graph passes through the given points. (-2,-3), (1,12),(-3,12)

OpenStudy (lalaly):

plug the points in the equation first equation -3=a(-2)^2+b(-2)+c -3=4a-2b+c second equation 12=a(1)^2+b(1)+c 12=a+b+c third equation 12=a(-3)^2+b(-3)+c 12=9a-3b+c now solve for the unkowns